# Real Bers Embedding on the Line: Fisher-Rao Linearization, Schwarzian Curvature, and Scattering Coordinates

### v1

@@ Abstract.
We develop a real-analytic counterpart of the Bers embedding for the Fréchet Lie group $\operatorname{Diff}^{-\infty}(\mathbb{R})$ of decay-controlled diffeomorphisms of the line, and establish its connection to $L^{p}$ Fisher-Rao geometry on densities. For $p\in[1,\infty)$, the $p$-root map $\varphi\mapsto p(\varphi^{\prime 1/p}-1)$ isometrically linearizes the homogeneous $\dot{W}^{1,p}$ Finsler metric on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, yielding explicit geodesics and a canonical flat connection whose Eulerian geodesic equation is the generalized Hunter-Saxton equation; for $p=\infty$, logarithmic coordinates $\varphi\mapsto\log\varphi^{\prime}$ provide a global isometry and the Schwarzian derivative emerges as the projective curvature. We construct a real Bers map $\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ via this Schwarzian, prove it is a Fréchet-smooth injective immersion whose linearization admits a tame right inverse given by an explicit Volterra operator, and characterize its image through Sturm-Liouville spectral theory. We introduce an $L^{p}$-Schwarzian family $S_{p}$ that interpolates between affine and projective cocycles, establish full asymptotic expansions as $p\to\infty$ in Fréchet and Orlicz-Sobolev scales, and extend the Bers embedding to Orlicz diffeomorphism groups. Through the Jacobian correspondence, these structures transfer to a manifold of densities asymptotic to Lebesgue measure, where the nonlinear Eulerian transport reduces to a pointwise Riccati law and the Schwarzian becomes the score curvature governing Fisher information. The compact-manifold $L^{p}$ Fisher-Rao linearization of Bauer, Bruveris, Harms, and Michor is recalled as a guiding framework.
@@
Schwarzian derivative, Bers embedding, Fisher-Rao metric, diffeomorphism groups, Orlicz spaces
@@ 1. Introduction
is a classical object in complex analysis and projective geometry, encoding the failure of a diffeomorphism to be a Möbius transformation. Its appearance in diverse contexts ranging from univalent function theory and Teichmüller spaces to conformal field theory and one-dimensional dynamics, which reflects a fundamental principle, that is, the Schwarzian is the unique differential invariant of third order that vanishes precisely on projective transformations and satisfies the chain rule cocycle
@@
These threads are linked by a single mechanism: the $p$-root map $\varphi\mapsto p(\varphi^{\prime 1/p}-1)$ isometrically flattens the right-invariant Finsler geometry on diffeomorphisms and, in the limit $p\to\infty$, produces logarithmic coordinates in which the Schwarzian takes the potential form $S(\varphi)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$, $f=\log(\varphi^{\prime})$. The real Bers embedding, the Sturm-Liouville spectral characterization of its image, the scattering transform, the Virasoro-Bott central extension, the $L^{p}$-Schwarzian interpolation, the Orlicz-Sobolev generalization, the Jacobian correspondence to densities, the forced sign-change theorems for projective curvature, and the sharp functional inequalities all follow from this single organizing principle.
Broader programme. This is the first of a two-part work. The one-dimensional scalar theory developed here serves as both the conceptual template and the technical foundation for Part II [21], which constructs a Koopman-equivariant Orlicz-Sobolev statistical bundle over a general non-singular dynamical system. There the Schwarzian becomes a Koopman spectral potential, the Bers embedding becomes a statistical bundle projection, and the scattering transform becomes an EDMD-type spectral decomposition. The blow-up analysis and spine decomposition of Part I supply the explicit one-dimensional inputs for the WKB expansions and Hamilton-Jacobi theory on the infinite-dimensional path space in Part II. The two parts together aim to provide a geometric and spectral framework in which the interplay between diffeomorphism groups, projective cocycles, statistical estimation, and branching process asymptotics can be treated in a unified manner.
@@ A.  Isometric linearization, flat connections, and geodesic dynamics
whose $x$-derivative gives the generalized Hunter-Saxton equation [16]. For $p=2$ this is the classical Hunter-Saxton equation. As $p\to\infty$, the quadratic nonlinearity vanishes and geodesics become solutions of the pure transport equation $\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})=0$. The connections $\nabla^{(p)}$ admit a two-parameter deformation $\nabla^{(p,\alpha)}$ by a projective tensor $\mathcal{C}(u,a)=u_{xx}a-ua_{xx}$; the resulting curvature involves the Lie bracket structure of vector fields through the commutator $\partial_{x}^{2},\mathcal{L}_{[u,v]}]$.
The variational formulation is equally clean: we prove an Euler-Lagrange characterization of geodesics (Theorem 2.20), establish the Hamiltonian duality through a Legendre transform on $L^{p}$-$L^{q}$ pairs (Theorem 2.21), and show that in the linearizing chart $\Phi_{p}$ the value function solves a Hamilton-Jacobi equation on the Banach space $L^{p}(\mathbb{R})$ with an explicit closed-form expression (Proposition 2.22). The family $\{\Phi_{p}\}_{p\in[1,\infty]}$ admits a full asymptotic expansion in Fréchet seminorms as $p\to\infty$ (Proposition 2.23).
B.  Blow-up analysis and large deviations
The Lagrangian strain $V(t,\xi)=u_{x}(t,\varphi(t,\xi))$ satisfies the autonomous Riccati ordinary differential equation
$\frac{\mathrm{d}V}{\mathrm{d}t}+\frac{1}{p}V^{2}=0$
at each label $\xi$, yielding the closed-form solution $V(t,\xi)=V_{0}(\xi)/(1+\frac{t}{p}V_{0}(\xi))$ and the Jacobian evolution $\partial_{\xi}\varphi(t,\xi)=\varphi_{0}^{\prime}(\xi)(1+\frac{t}{p}V_{0}(\xi))^{p}$. From this we derive a sharp blow-up criterion (Theorem 3.3): the geodesic loses smoothness at time $T^{*}=-p/\inf(v_{0})$ if and only if $\inf(v_{0})<0$, where the strain develops a $|x-x^{*}|^{-1}$ singularity with an explicit self-similar profile (Theorem 3.5).
Beyond blow-up, we prove global existence and uniqueness of dissipative weak solutions for the generalized Hunter-Saxton equation (Theorem 3.7) extending the framework of [9] from $p=2$ to general $p$. Geometrically, the dissipative continuation corresponds to the affine geodesic in $L^{p}(\mathbb{R})$ exiting the positivity constraint $\{f>-p\}$-the nonlinearity of the diffeomorphism group is encoded entirely in this half-space condition.
The blow-up admits a probabilistic interpretation through a branching factor and local Malthusian parameter (Definitions 3.8). We establish a small-noise large deviation principle for the catastrophe time $T^{*}$ under Gaussian perturbation of the initial strain (Theorem 3.18), with a quadratic rate function $J(T)\sim\frac{m^{4}}{2p^{2}\kappa}(T-T^{*})^{2}$ governed by the Cameron-Martin geometry. A spine decomposition (Proposition 3.20, Theorem 3.21) shows that the unique minimal-cost perturbation achieving a prescribed shift in blow-up time is proportional to the Green function $\frac{1}{2}e^{-|x-x^{*}|}$, the Riesz representer of point evaluation in $H^{1}(\mathbb{R})$.
C.  The Schwarzian as projective curvature
@@
Theorem 1.2 ($L^{p}$-Schwarzian, Propositions 4.20 and 4.24).
@@
The Sturm-Liouville equation $y^{\prime\prime}+qy=0$ provides a canonical differential-geometric framework for the Schwarzian, and we develop a semiclassical interpretation in which the WKB expansion of solutions reproduces the classical and $L^{p}$ Schwarzians as successive curvature corrections (Section 4.7).
D.  Central extensions and cocycle geometry
@@
A new ingredient is the cumulant transgression: although the Bott cocycle is not a group coboundary on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, it becomes an explicit groupoid coboundary on the action groupoid $\operatorname{Diff}^{-\infty}(\mathbb{R})\ltimes|\Omega|_{\infty,1}(\mathbb{R})$ once one allows dependence on a density through its score (Section 5, Subsection 3.6). This yields a canonical transgression functional whose coboundary is the Bott phase, providing a categorical absorption of the projective multiplier into a central extension of Koopman transport. We formalize this through a $U(1)$-banded gerbe on the density quotient.
E.  The real Bers embedding
@@
We also construct local Bers coordinates centered at arbitrary basepoints, compute their transition maps as Schwarzian-shifted affine connections (Section 6.8), and establish a spectral transport formula that conjugates Schrödinger operators along paths in the Bers image (Section 6.9).
F.  Scattering coordinates and nonlinear spectral analysis
@@
We develop the scattering theory in detail. Jost solutions and their Volterra integral equations are constructed for potentials in $W^{\infty,1}(\mathbb{R})$, yielding Schwartz-class scattering data (Section 7). We introduce a relative spectral zeta function $\zeta_{\mu}(z)$ for the Bers Schrödinger operator and prove that its residue at $z=-\frac{1}{2}$ is a multiple of the Fisher information (Proposition 7.8). The Bers-Mercer kernel $K_{q,\rho}$, obtained by integrating the Green function of $H_{q}$ against an admissible spectral weight, is shown to be positive definite, giving rise to a reproducing kernel Hilbert space $\mathcal{H}_{q,\rho}$ that provides the canonical arena for finite-rank spectral approximation (Section 7.6). The Marchenko reconstruction is developed in Section 7.7.
G.  Extension to Orlicz-Sobolev scales
The entire theory extends to Orlicz-Sobolev scales. For an admissible Young function $\Phi$ satisfying a $\Delta_{2}$ condition, we construct the Orlicz diffeomorphism group $\operatorname{Diff}^{-\Phi}(\mathbb{R})$ in logarithmic coordinates and prove it carries a Fréchet Lie group structure (Section 8). The construction proceeds through Orlicz fiber bundles over the density manifold $|\Omega|_{\infty,1}(\mathbb{R})$, equipped with the Luxemburg norm fiberwise. We develop the Orlicz-Young duality on fibers (Section 8.5), prove that Koopman operators act as bounded bundle morphisms on the Orlicz fibers (Section 8.6), and establish that the Bers embedding extends to a Fréchet smooth injective immersion $\beta_{\Phi}\colon\operatorname{Diff}^{-\Phi}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\to W^{\infty,\Phi}(\mathbb{R})$. The convergence $S_{p}\to S$ is proved to be stable in Orlicz-Sobolev norms (Section 5.8).
H.  Density geometry: projective curvature and forced sign-change
@@
We develop the Koopman operator theory on density space, constructing a statistical Hilbert bundle with a unitary half-density representation and a half-density generator whose drift is the score (Section 9.5). The projective covariance of the Koopman action produces the Schwarzian cocycle (Theorem 9.19). A Feynman-Kac representation identifies $\beta(\mu)/2$ as the potential of a branching-killing diffusion (Theorem 9.38), and the forced sign-change theorem translates into the statement that every nontrivial density has both a genuine branching region and a genuine killing region. The spine change of measure identifies the score $s_{\mu}$ as (twice) the drift of the spine process (Theorem 9.40).
I.  Sharp functional inequalities
@@
The deepest results in this section are trace identities that express the geometric integrals of the Bers potential in spectral variables. The first trace identity, which equates $\int q\,dx$ with a spectral entropy of the reflection coefficient (Theorem 10.10), is given by
@@ 1.2. Relation to prior work
The Fréchet Lie group structure on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and the existence of right-invariant Sobolev metrics were established by Bauer, Bruveris, and Michor [4]. The classical Bers embedding in Teichmüller theory is treated in Lehto [22]. The Schwarzian derivative and its role in projective geometry are classical; see Ovsienko and Tabachnikov [25] for a modern treatment and Fuks [14] for the cohomological aspects. The Hunter-Saxton equation as a geodesic equation on diffeomorphism groups was identified by Khesin and Misiołek [18]; the original derivation of the equations is in [16]. The geometry of infinite-dimensional groups is treated comprehensively in Khesin and Wendt [19]. The one-dimensional inverse scattering theory on the line is due to Deift and Trubowitz [10]; see also Beals, Deift, and Tomei [7] and Faddeev and Takhtajan [12]. The Miura transformation connecting KdV and mKdV equations originates in [23]. The Bott-Thurston cocycle and its role in the theory of characteristic classes of diffeomorphism groups appears in Bott [8] and the Virasoro extension in Pressley and Segal [26]; the cohomological framework is developed in Feigin and Fuchs [13] and Fuks [14]. The integration of the Lie algebra cocycle to a group cocycle follows the general pattern of [15]. We verify the details in the present decay-controlled setting. Orlicz space theory follows Rao and Ren [27] and the Sobolev embedding theorems are from Adams and Fournier [1].
@@ 1.3. Organization
The paper is organized as follows.
Section 2 establishes the isometric linearization of $\operatorname{Diff}^{-\infty}(\mathbb{R})$, derives the flat connection, geodesics, parallel transport, the Hamiltonian formulation with Hamilton-Jacobi equation, the two-parameter family of connections with its torsion and curvature, and the Fréchet asymptotics as $p\to\infty$.
Section 3 develops the complete blow-up theory for the generalized Hunter-Saxton equation: the Lagrangian Riccati structure, sharp blow-up criterion and rate, global existence and uniqueness of dissipative weak solutions, the probabilistic branching-factor interpretation, and the large deviation principle for the catastrophe time with its spine decomposition.
Section 4 develops the Schwarzian derivative as a projective potential in logarithmic coordinates, introduces the score curvature and its Fisher-Rao identities including the Bhattacharyya-type bound, constructs the projective connection bundle, introduces the $L^{p}$-Schwarzian family with its cocycle identity and $p\to\infty$ asymptotics, and gives the Sturm-Liouville and semiclassical interpretations.
Section 5 treats central extensions and cocycle geometry: the Gelfand-Fuchs cocycle and its nontriviality, the Virasoro-Bott group extension, $L^{p}$ cocycle rigidity, the Bott-Thurston cocycle in $p$-root coordinates, the gerbe structure on the density quotient, and the cumulant transgression producing Koopman central extensions.
Section 6 constructs the real Bers embedding, proves injectivity and the tame right inverse theorem, establishes the Miura factorization, the spectral characterization of the image, the explicit inverse construction, local Bers coordinates with their transition maps, and spectral transport along paths.
Section 7 develops the scattering theory for Bers potentials: Jost solutions, the absence of bound states, the scattering transform as a nonlinear Fourier map, spectral zeta functions and regularized determinants, Mercer kernels and reproducing kernel Hilbert spaces, and the Marchenko reconstruction.
Section 8 extends the theory to Orlicz-Sobolev scales: Orlicz-Sobolev spaces, admissible Young functions, Orlicz diffeomorphism groups, Orlicz fiber bundles and their Koopman theory, Orlicz-Young duality, the Orlicz Bers embedding, and the Orlicz stability of $S_{p}\to S$.
Section 9 develops the density-side theory: the Jacobian isomorphism, score functions and projective curvature, the linearization intertwining, the forced sign-change theorems with their inflection-point and Fisher non-control consequences, the Koopman operators and statistical Hilbert bundle with unitary half-density representation, the path space Fisher-Rao action, and the branching diffusion viewpoint with Feynman-Kac and spine constructions.
Section 10 establishes the sharp functional inequalities: critical Hardy theory and resonance rigidity, Hardy-Poincaré coercivity on finite windows, $p$-Fisher energy theory, and the trace hierarchy expressing Fisher energies as spectral entropies of the reflection coefficient.
Section 11 recalls the $L^{p}$ Fisher-Rao framework on compact manifolds and the BBM classification theorem, serving as context for the noncompact extensions developed in the body of the paper.
1.4. Relation to Part II
This paper is the first of a two-part work. Part II [21] constructs the Koopman-equivariant Orlicz-Sobolev statistical bundle over a general measure-preserving or non-singular dynamical system, where the one-dimensional structures developed here lift to an infinite-dimensional fiber bundle. Specifically, the Luxemburg-Finsler geometry from Section 8, the Koopman action on Orlicz fibers from Section 8.6, and the cumulant transgression from Section 5 become the building blocks for a bundle of Orlicz function spaces on which non-singular dynamics act as bundle automorphisms preserving the Finsler structure.
In that setting, the Bers-Mercer kernels of Section 7.6 yield invariant reproducing kernels that enable spectral reductions, and the blow-up analysis and spine decomposition of Section 3 supply the one-dimensional WKB inputs for a semiclassical Hamilton-Jacobi problem on the path space of multitype branching processes. This produces WKB expansions with Berry-phase corrections, where the phase is controlled by the Bott-Thurston cocycle (Section 5), and an explicit spine action functional whose critical points recover the large deviation rate. The trace identities of Section 10 provide the spectral entropy variables that serve as the correct action-angle coordinates for the infinite-dimensional problem.

### v2

# The Real Bers Embedding on the Line: Fisher-Rao Linearization, Schwarzian Curvature, and Scattering Coordinates

@@ Abstract.
We develop a real-analytic counterpart of the Bers embedding for the Fréchet Lie group $\operatorname{Diff}^{-\infty}(\mathbb{R})$ of decay-controlled diffeomorphisms of the line. For $p\in[1,\infty)$, the $p$-root map $\varphi\mapsto p(\varphi^{\prime 1/p}-1)$ isometrically linearizes the homogeneous $\dot{W}^{1,p}$ Finsler metric on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, yielding explicit geodesics and a canonical flat connection whose Eulerian geodesic equation is the generalized Hunter-Saxton equation; for $p=\infty$, logarithmic coordinates provide a global isometry and the Schwarzian derivative emerges as the projective curvature. We construct a real Bers map $\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ via this Schwarzian, prove it is a Fréchet-smooth injective immersion whose linearization admits a tame right inverse given by an explicit Volterra operator, and characterize its image through Sturm-Liouville spectral theory and one-dimensional scattering. Through the Jacobian correspondence, these structures transfer to a manifold of densities asymptotic to Lebesgue measure, where the Schwarzian becomes the score curvature governing Fisher information. We prove that for every nontrivial density, the projective curvature necessarily takes both signs (forced sign-change), while Fisher information alone does not control the positive projective curvature. Trace identities express Fisher information and its higher analogues as spectral entropies of the scattering reflection coefficient.
@@
Schwarzian derivative, Bers embedding, Fisher-Rao metric, diffeomorphism groups, scattering theory
@@ 1. Introduction
is a classical object in complex analysis and projective geometry, encoding the failure of a diffeomorphism to be a Möbius transformation. Its appearance in diverse contexts, from univalent function theory and Teichmüller spaces to conformal field theory and one-dimensional dynamics, reflects a fundamental principle: the Schwarzian is the unique differential invariant of third order that vanishes precisely on projective transformations and satisfies the chain rule cocycle
@@
These threads are linked by a single mechanism: the $p$-root map $\varphi\mapsto p(\varphi^{\prime 1/p}-1)$ isometrically flattens the right-invariant Finsler geometry on diffeomorphisms and, in the limit $p\to\infty$, produces logarithmic coordinates in which the Schwarzian takes the potential form $S(\varphi)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$, $f=\log(\varphi^{\prime})$. The real Bers embedding, the Sturm-Liouville spectral characterization of its image, the scattering transform, the Virasoro-Bott central extension, the $L^{p}$-Schwarzian interpolation, the Jacobian correspondence to densities, the forced sign-change theorems for projective curvature, and the sharp functional inequalities all follow from this single organizing principle.
@@ A.  Isometric linearization, flat connections, and geodesic dynamics
whose $x$-derivative gives the generalized Hunter-Saxton equation. For $p=2$ this is the classical Hunter-Saxton equation. As $p\to\infty$, the quadratic nonlinearity vanishes and geodesics become solutions of the pure transport equation $\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})=0$. The connections $\nabla^{(p)}$ admit a two-parameter deformation $\nabla^{(p,\alpha)}$ by a projective tensor $\mathcal{C}(u,a)=u_{xx}a-ua_{xx}$; the resulting curvature involves the Lie bracket structure of vector fields through the commutator $\partial_{x}^{2},\mathcal{L}_{[u,v]}]$.
The family $\{\Phi_{p}\}_{p\in[1,\infty]}$ admits a full asymptotic expansion in Fréchet seminorms as $p\to\infty$ (Proposition 2.16).
B.  The Schwarzian as projective curvature
@@
Theorem 1.2 ($L^{p}$-Schwarzian, Propositions 3.20 and 3.24).
@@
The Sturm-Liouville equation $y^{\prime\prime}+qy=0$ provides a canonical differential-geometric framework for the Schwarzian.
C.  Central extensions and cocycle geometry
@@
D.  The real Bers embedding
@@
E.  Scattering coordinates and nonlinear spectral analysis
@@
We develop the scattering theory in detail. Jost solutions and their Volterra integral equations are constructed for potentials in $W^{\infty,1}(\mathbb{R})$, yielding Schwartz-class scattering data (Section 6).
F.  Density geometry: projective curvature and forced sign-change
@@
G.  Sharp functional inequalities
@@
The main analytic results in this section are trace identities that express the geometric integrals of the Bers potential in spectral variables. The first trace identity equates $\int q\,dx$ with a spectral entropy of the reflection coefficient (Theorem 8.7):
@@ 1.2. Relation to prior work
The Fréchet Lie group structure on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and the existence of right-invariant Sobolev metrics were established by Bauer, Bruveris, and Michor [2]. The classical Bers embedding in Teichmüller theory is treated in Lehto [14]. The Schwarzian derivative and its role in projective geometry are classical; see Ovsienko and Tabachnikov [16] for a modern treatment and Fuks [10] for the cohomological aspects. The Hunter-Saxton equation as a geodesic equation on diffeomorphism groups was identified by Khesin and Misiołek [11]; the geometry of infinite-dimensional groups is treated comprehensively in Khesin and Wendt [12]. The one-dimensional inverse scattering theory on the line is due to Deift and Trubowitz [7]; see also Beals, Deift, and Tomei [5] and Faddeev and Takhtajan [8]. The Miura transformation connecting KdV and mKdV equations originates in [15]. The Bott-Thurston cocycle and its role in the theory of characteristic classes of diffeomorphism groups appears in Bott [6] and the Virasoro extension in Pressley and Segal [17]; the cohomological framework is developed in Feigin and Fuchs [9] and Fuks [10]. Orlicz space theory follows Rao and Ren [18] and the Sobolev embedding theorems are from Adams and Fournier [1].
@@ 1.3. Organization
The paper is organized as follows. Section 2 establishes the isometric linearization of $\operatorname{Diff}^{-\infty}(\mathbb{R})$, derives the flat connection, geodesics, and the Fréchet asymptotics as $p\to\infty$. Section 3 develops the Schwarzian derivative as a projective potential, introduces the score curvature and its Fisher-Rao identities, constructs the $L^{p}$-Schwarzian family, and gives the Sturm-Liouville interpretation. Section 4 treats the Gelfand-Fuchs cocycle, the Virasoro-Bott group extension, and $L^{p}$ cocycle rigidity. Section 5 constructs the real Bers embedding, proves injectivity and the tame right inverse theorem, establishes the Miura factorization, and gives the spectral characterization of the image. Section 6 develops the scattering theory for Bers potentials: Jost solutions, the absence of bound states, and the scattering transform as a nonlinear Fourier map. Section 7 develops the density-side theory: the Jacobian isomorphism, score functions and projective curvature, the forced sign-change theorems with their inflection-point and Fisher non-control consequences. Section 8 establishes the sharp functional inequalities: critical Hardy theory, and the trace hierarchy expressing Fisher energies as spectral entropies of the reflection coefficient. An appendix recalls the $L^{p}$ Fisher-Rao framework on compact manifolds.
Broader programme. This paper is the first of a two-part work. The blow-up analysis for the generalized Hunter-Saxton equation (large deviations and spine decomposition), the extension to Orlicz-Sobolev scales, gerbe structures on the density quotient, and the Koopman-equivariant statistical bundle over general non-singular dynamical systems will appear in a companion paper and in Part II [13].